OpenStudy (anonymous):

write an equation of the circle with center (-2,-3) and radius 8 i did x^2+4+y^2+9=64

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

why not the standard quation is (x-h)^2+(y-k)^2=r^2 and i plugged in the values

OpenStudy (anonymous):

circle equation \[(x-x_o)^2+(y-y_o)^2=r^2\] if i remembered correctly anyways \[(x+2)^2 \ne x^2+4\]

OpenStudy (anonymous):

you used the right equation, you just "foil"ed incorrectly

OpenStudy (anonymous):

ohhh i get it (x+2)(x+2)

OpenStudy (anonymous):

 yup